OpenStudy (anonymous):

how does 2a-2b/5a-5b become 2/5 ?

OpenStudy (anonymous):

\[\frac{ 2a-2b }{ 5a-5b}=\frac{ 2(a-b) }{ 5(a-b)} , a \neq b \]

OpenStudy (anonymous):

\[\frac{ a-b }{ a-b}=1 ,a \neq b\]